# Chain-of-Frames: Advancing Video Understanding in Multimodal LLMs via Frame-Aware Reasoning

### v1

@@ A.2 Video benchmarks
Video-MME. Video-MME (Fu et al., 2024) offers a diverse range of video types, covering six primary visual domains and 30 subfields to support broad scenario generalizability. It also introduces variation in temporal length, including short (under 2 minutes), medium (4-15 minutes), and long (30-60 minutes) videos.
MVBench. Li et al. (2023b) presents a comprehensive benchmark for multimodal video understanding, encompassing 20 challenging tasks that require more than single-frame analysis. It is specifically designed to evaluate a model’s ability to understand temporal dynamics across video sequences.
VSI-Bench. This benchmark (Yang et al., 2024) is designed to quantitatively assess the visual-spatial intelligence of multimodal large language models. Built from over 5,000 high-quality question-answer pairs across 288 real-world indoor videos, VSI-Bench spans diverse environments such as homes, offices, and industrial spaces. The benchmark covers eight tasks: object count, relative distance, relative direction, route planning, object size estimation, room size estimation, absolute distance estimation, and appearance order. Out of these tasks included in this benchmark, relative distance, appearance order, relative directory, and route planning come with multiple-choice questions while the other four require an open-ended quantitative answer. To better evaluate the proximity of the model’s prediction with the correct answer, Yang et al. (2024) proposes using mean relative accuracy ($\mathcal{MRA}$). Given a model’s prediction $\hat{y}$ and ground truth $y$, relative accuracy is calculated by:
@@
VidHal. To evaluate video-based hallucinations in video LLMs, we use VidHal (Choong et al., 2024), a multiple-choice benchmark that features video instances drawn from public video understanding datasets, covering a diverse array of temporal concepts and aspects such as entity actions and event sequences.
EventHallusion. Zhang et al. (2024a) introduce EventHallusion, from which we use the binary-choice questions designed to systematically assess event-related hallucinations in video LLMs. From a hallucination attribution standpoint, it is specifically curated to evaluate a model’s susceptibility to language priors and vision-language correlation biases.
@@ A.3 Chain-of-Frames model
For InternVL2.5-4B, we fully fine-tune both the LLM and the projection modules, keeping the vision encoder frozen. In contrast, for InternVL3-8B, we adopt LoRA-based fine-tuning (Hu et al., 2022) to reduce memory consumption. All other training configurations remain consistent across both models. Training is conducted on a single H100 node equipped with 4 GPUs, using a learning rate of $2\times 10^{-6}$, a batch size of 2, and a single epoch.
@@ Appendix B Additional Experiments
Comparison to additional baselines. We could not include the original models from other works (Wang et al., 2024b; Han et al., 2024; Fei et al., 2024; Hu et al., 2025b) in the main evaluation in Table 2 because they are not accessible to the public, and they do not report results on the five benchmarks we use. However, we provide comparative analyses to M-LLM (Hu et al., 2025b) and Video-of-Thought (Fei et al., 2024) on Video-MME and NextQA (Xiao et al., 2021) using the results they report. In Table 3, we see that our CoF-based models, including CoF-InternVL2.5-4B which has fewer parameters than the competitors, outperform these baselines on both benchmarks. Moreover, compared to the most recent model, M-LLM, our method achieves a greater improvement (4.9% vs 0.8%) on NextQA despite starting from a stronger base model.
Table 3: Comparison to additional baselines. Since the models from Hu et al. (2025b) and Fei et al. (2024) are not publicly available, we compare the results they report on Video-MME and NextQA to those of our CoF models. We also report the video LLM each model is based on and its performance, as well as the improvement of each approach over the original model. Both our CoF models outperform the existing approaches on the two benchmarks.
Backbone Model Video-MME NextQA
Video-LLaVA-7B Original 39.9 66.3
Video-of-Thought (Fei et al., 2024) - 76.0  9.7$\uparrow$
Qwen2-VL-7B Original 58.1 77.6
M-LLM (Hu et al., 2025b) 58.7  0.6$\uparrow$ 78.4  0.8$\uparrow$
InternVL2.5-4B Original 54.9 75.3
SFT with CoF (ours) 59.7  4.8$\uparrow$ 79.6  4.3$\uparrow$
InternVL3-8B Original 66.5 82.4
SFT with CoF (ours) 75.38.8$\uparrow$ 87.34.9$\uparrow$
Effect of CoT prompting. An extended version of Table 1 is presented in Table 4. For all baselines, we report results using two prompting strategies, either standard (indicated by $\bm{\star}$) or chain-of-thought (indicated by $\clubsuit$, the prompt is shown in Fig. 10). For our SFT with CoF models, we always use CoT prompting. When considering InternVL2.5-4B, CoT prompting alone improves the accuracy of the original models on four out of five benchmarks compared to the original model. However, this improvement does not hold for the SFT with QA only variant: we hypothesize that fine-tuning solely on QA data negatively impacts the reasoning capabilities of the baseline model. On the other hand, incorporating reasoning traces into the training data (SFT with CoT) generally enhances the model’s reasoning capabilities, and using CoT prompting is beneficial except for the EventHallusion benchmark. CoT prompting improves the results also for the original InternVL3-8B on most benchmarks. Finally, our models (SFT with CoF) outperform the baseline across all benchmarks.
Table 4: Effect of CoT prompting. For all baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix CoT prompting for our CoF models.
@@
SFT with CoF (ours) $\clubsuit$ 51.3 75.3 90.4 79.5 78.7
Detailed results over benchmark splits. For completeness, we report the fine-grained results over the various splits of VSI-Bench (Table 5), Video-MME (Table 6), MVBench (Table 7), and EventHallusion (Table 8). Moreover, for the baseline models, we report results using two prompting strategies, i.e., standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$).
Detailed results of CoF reasoning at inference time. In Fig. 11, we show statistics of how many frames are referenced in the reasoning traces generated by CoF-InternVL3-8B. To complement Fig. 7, we report the frequency for each benchmark separately. We see that the number of frames mentioned varies across benchmarks, e.g., the cases where no frames are referenced significantly decreases on the hallucination benchmarks VidHal and EventHallusion. This suggests that our CoF models learn to modulate the reasoning traces and the frame references depending on the task.
@@
Table 5: Detailed results on the VSI-Bench benchmark. For the baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
@@
Table 6: Detailed results on the Video-MME benchmark. For the baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
@@
SFT with CoF (ours) $\clubsuit$ 80.9 74.4 70.4 75.3
Table 7: Detailed results on the MVBench benchmark. For the baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.
@@
Original $\bm{\star}$ 90.0 42.0 44.5 83.0 82.5 75.5 78.0 38.5 45.0 98.0
$\clubsuit$ 77.5 45.0 42.5 87.5 97.9 80.5 89.0 33.5 54.0 99.0
SFT with CoF (ours) $\clubsuit$ 96.5 95.5 58.0 99.5 100.0 99.0 77.5 79.5 87.5 96.5
@@
Original $\bm{\star}$ 60.5 89.0 97.0 85.5 39.5 92.5 69.0 81.0 72.0
$\clubsuit$ 70.5 91.0 97.5 87.5 41.5 92.5 77.0 73.5 74.3
SFT with CoF (ours) $\clubsuit$ 85.0 89.0 90.4 96.5 95.5 94.0 99.0 88.0 90.4
Table 8: Detailed results on the VidHal and EventHallusion benchmarks. For the baselines, we report results using two prompting strategies, i.e. standard (indicated by $\bm{\star}$) and chain-of-thought (indicated by $\clubsuit$), while we fix chain-of-thoughts prompting for our CoF models.

### v2

@@ A.2 Video benchmarks
Video-MME.
Video-MME [10] offers a diverse range of video types, covering six primary visual domains and 30 subfields to support broad scenario generalizability. It also introduces variation in temporal length, including short (under 2 minutes), medium (4-15 minutes), and long (30-60 minutes) videos.
MVBench.
Li et al. [26] presents a comprehensive benchmark for multimodal video understanding, encompassing 20 challenging tasks that require more than single-frame analysis. It is specifically designed to evaluate a model’s ability to understand temporal dynamics across video sequences.
VSI-Bench.
This benchmark [38] is designed to quantitatively assess the visual-spatial intelligence of multimodal large language models. Built from over 5,000 high-quality question-answer pairs across 288 real-world indoor videos, VSI-Bench spans diverse environments such as homes, offices, and industrial spaces. The benchmark covers eight tasks: object count, relative distance, relative direction, route planning, object size estimation, room size estimation, absolute distance estimation, and appearance order. Out of these tasks included in this benchmark, relative distance, appearance order, relative directory, and route planning come with multiple-choice questions while the other four require an open-ended quantitative answer. To better evaluate the proximity of the model’s prediction with the correct answer, [38] proposes using mean relative accuracy ($\mathcal{MRA}$). Given a model’s prediction $\hat{y}$ and ground truth $y$, relative accuracy is calculated by:
@@
VidHal.
To evaluate video-based hallucinations in video LLMs, we use VidHal [8], a multiple-choice benchmark that features video instances drawn from public video understanding datasets, covering a diverse array of temporal concepts and aspects such as entity actions and event sequences.
EventHallusion.
Zhang et al. [40] introduce EventHallusion, from which we use the binary-choice questions designed to systematically assess event-related hallucinations in video LLMs. From a hallucination attribution standpoint, it is specifically curated to evaluate a model’s susceptibility to language priors and vision-language correlation biases.
@@ A.3 Chain-of-Frames model
CoF-InternVL.
For InternVL2.5-4B, we fully fine-tune both the LLM and the projection modules, keeping the vision encoder frozen. In contrast, for InternVL3-8B, we adopt LoRA-based fine-tuning [15] to reduce memory consumption. All other training configurations remain consistent across both models. Training is conducted on a single H100 node equipped with 4 GPUs, using a learning rate of $2\times 10^{-6}$, a batch size of 2, and a single epoch.
CoF-Phi-3.5-Vision-4B.
To test the generalizability of our approach beyond the InternVL family, we employ Phi-3.5-Vision-4B [1], a mobile-scale multimodal LLM that demonstrates strong performance in language reasoning, as a third baseline model.
Phi-3.5-Vision-4B uses distinct, indexed image placeholder tokens such as <imagei-i>, instead of repeating a generic <image> token, because the model must uniquely align each image embedding with a specific position in the text sequence. Each placeholder is a separate token in the tokenizer, allowing the model to reliably map image embeddings to their corresponding locations and to correctly understand references to the images.
Phi-3.5-Vision’s capability to understand references to the images makes it a good candidate for our base model. Similar to InternVL2.5-4B, we fully fine-tune both the LLM and the projection modules, keeping the vision encoder frozen. The other training configurations remain consistent across all baselines.
@@ Appendix B Additional Experiments
CoF-Phi-3.5-Vision-4B results.
Tab. 4 compares the baseline Phi-3.5-Vision-4B model with its CoF-enhanced variant. Across all benchmarks, CoF-Phi-3.5-Vision-4B consistently surpasses the corresponding base model. The largest improvements appear on the VSI-Bench benchmark, where CoF-Phi-3.5-Vision-4B achieves better performance compared to LLaVA-OneVision-7B and Qwen2-VL-7B (Tab. 1) despite using substantially fewer parameters. These results demonstrate that our chain-of-frames approach generalizes effectively across architectures and model families.
Effect of CoT prompting.
[…]
@@
SFT with CoF (ours) $\clubsuit$ 51.3 73.7 77.1 79.5 78.7
Phi-3.5-Vision-4B
[…]
$\clubsuit$ 26.4 39.0 46.6 43.1 51.3
SFT with CoF (ours) $\clubsuit$ 32.8 52.6 51.8 55.5 59.2
[…]
Detailed results over benchmark splits.
[…]
Detailed results of CoF reasoning at inference time.
In Fig. 12, we show statistics of how many frames are referenced in the reasoning traces generated by CoF-InternVL3-8B. To complement Fig. 7, we report the frequency for each benchmark separately. We see that the number of frames mentioned varies across benchmarks, e.g., the cases where no frames are referenced significantly decreases on the hallucination benchmarks VidHal and EventHallusion. This suggests that our CoF models learn to modulate the reasoning traces and the frame references depending on the task.
@@
Phi-3.5-Vision-4B
[…]
$\clubsuit$ 27.3 24.1 32.6 22.6 20.5 20.9 35.2 28.3 26.4
SFT with CoF (ours) $\clubsuit$ 31.5 27.9 37.9 26.3 33.5 22.7 47.3 35.1 32.8
[…]
@@
SFT with CoF (ours) $\clubsuit$ 79.3 71.7 70.0 73.7
Phi-3.5-Vision-4B
[…]
$\clubsuit$ 46.0 34.9 36.1 39.0
SFT with CoF (ours) $\clubsuit$ 63.1 51.9 42.7 52.6
[…]
@@
[…]
$\clubsuit$ 77.5 45.0 42.5 87.5 85.1 80.5 89.0 33.5 47.0 99.0
SFT with CoF (ours) $\clubsuit$ 96.5 50.5 49.5 89.5 91.0 91.5 77.5 45.5 59.5 96.5
Phi-3.5-Vision-4B
[…]
$\clubsuit$ 61.0 49.5 27.5 43.0 37.8 42.0 35.0 36.5 36.5 61.5
SFT with CoF (ours) $\clubsuit$ 68.0 51.5 38.5 49.5 42.02 49.5 39.5 44.0 38.5 68.0
@@
[…]
$\clubsuit$ 63.5 89.0 97.47 87.5 41.5 92.5 73.0 73.5 72.5
SFT with CoF (ours) $\clubsuit$ 67.0 91.0 90.4 90.5 45.5 94.0 78.0 84.0 77.1
Phi-3.5-Vision-4B
[…]
$\clubsuit$ 35.5 38.5 53.0 43.5 28.5 89.0 47.5 72.5 46.6
SFT with CoF (ours) $\clubsuit$ 36.5 45.5 61.1 46.5 33.5 90.0 53.3 76.5 51.8
[…]